OpenStudy (anonymous):

A simple pendulum is 5 m long. What is the period of small oscillations for this pendulum if it is located in an elevator (a) Standing still? (b) Moving upwards with constant velocity of 7.45 m/s? (c) Accelerating upward at 5 m/s2? (d) Accelerating downward at 5 m/s2?

OpenStudy (priyar):

if the oscillations are small, it will undergo SHM .. so time period(T) is given by: \[T = 2\pi \sqrt{\frac{ l }{ g }}\] where l-length and g- gravitational acceleration..

OpenStudy (priyar):

l is constant and is equal to 5m. So decide how g will change in each case..